### v1

# Multiplayer Nash Preference Optimization

Abstract
Reinforcement learning from human feedback (RLHF) has emerged as the standard paradigm for aligning large language models (LLMs) with human preferences. However, reward-based methods built on the Bradley–Terry assumption struggle to capture the non-transitive and heterogeneous nature of real-world preferences. To address this, recent studies have reframed alignment as a two-player Nash game, giving rise to Nash learning from human feedback (NLHF). While this perspective has inspired algorithms such as INPO, ONPO, and EGPO with strong theoretical and empirical guarantees, they remain fundamentally restricted to two-player interactions, creating a single-opponent bias that fails to capture the full complexity of realistic preference structures. In this work, we introduce Multiplayer Nash Preference Optimization (MNPO), a novel framework that generalizes NLHF to the multiplayer regime. It formulates alignment as an $n$-player game, where each policy competes against a population of opponents while being regularized toward a reference model. Our framework establishes well-defined Nash equilibria in multiplayer settings and extends the concept of duality gap to quantify approximation quality. We demonstrate that MNPO inherits the equilibrium guarantees of two-player methods while enabling richer competitive dynamics and improved coverage of diverse preference structures. Through comprehensive empirical evaluation, we show that MNPO consistently outperforms existing NLHF baselines on instruction-following benchmarks, achieving superior alignment quality under heterogeneous annotator conditions and mixed-policy evaluation scenarios. Together, these results establish MNPO as a principled and scalable framework for aligning LLMs with complex, non-transitive human preferences. Code is available at https://github.com/smiles724/MNPO.
1 Introduction
Large language models (LLMs) have achieved remarkable progress in instruction following and open-ended reasoning, primarily through reinforcement learning from human feedback (RLHF) (Christiano et al., 2017). In RLHF, annotators provide preference comparisons between model outputs, and these signals are used to align LLMs with human expectations. While reward-based RLHF pipelines built upon the Bradley-Terry (Bradley & Terry, 1952) model have enabled widely deployed systems (e.g., InstructGPT (Ouyang et al., 2022), Claude (Bai et al., 2022), and Gemini (Team et al., 2023)), they assume transitive preferences and scalar reward functions. Recent empirical studies in LLM alignment have revealed that human preferences often exhibit non-transitive patterns and heterogeneous structures that challenge these assumptions (Ethayarajh et al., 2024; Wu et al., 2024).
This limitation has motivated recent attempts to develop game-theoretic formulations of alignment. Rather than relying on scalar rewards that enforce transitivity constraints, recent work treats preference optimization as finding Nash equilibria in games defined by general preference oracles (Munos et al., 2023). In this Nash learning from human feedback (NLHF) paradigm, a well-aligned policy represents an optimal strategy that cannot be exploited by competing policies. At equilibrium, policies achieve a balanced win rate that reflects strategic optimality rather than mediocrity. Subsequent studies have explored this paradigm through various algorithmic lenses. For instance, Zhang et al. (2025b) leverages no-regret learning to approximate the Nash equilibrium via self-play. Zhang et al. (2025a) accelerates convergence using optimistic mirror descent. Zhou et al. (2025) establishes last-iterate convergence with extragradient updates. These approaches advance both theoretical guarantees and empirical stability compared to traditional RLHF.
Despite advances, the existing NLHF remains constrained to two-player settings where a single policy competes against one opponent. However, real-world preference alignment often involves diverse annotators, heterogeneous evaluation criteria, or mixtures of historical model checkpoints - contexts that are better modeled as multiplayer games (Freund & Schapire, 1999). Extending NLHF to the multiplayer regime raises new challenges: (i) how to formulate objectives that balance fairness and competitiveness among multiple policies, (ii) how to design algorithms that converge to multiplayer Nash equilibria, and (iii) how to ensure tractability when scaling to larger LLMs (Sokota et al., 2022).
We address these challenges by proposing Multiplayer Nash Preference Optimization (MNPO), a principled framework that generalizes two-player preference optimization to $n$-player games. In MNPO, each policy simultaneously competes against all other policies while being regularized toward a reference policy, creating a competitive equilibrium that balances performance against the population with adherence to a trusted baseline. Our contributions are threefold:
  • Theoretical Framework: We establish that MNPO admits natural equilibrium characterizations, including well-defined Nash policies and duality gaps that measure alignment quality. We prove that MNPO inherits the desirable convergence properties of existing two-player formulations while enabling richer equilibrium dynamics.
  • Algorithmic Innovation: We introduce time-dependent MNPO (TD-MNPO), where opponent sets evolve adaptively using weighted combinations of historical policies. This approach enables models to dynamically incorporate past knowledge while maintaining training stability.
  • Empirical Validation: Through comprehensive experiments on instruction-following and reasoning benchmarks, we demonstrate that MNPO consistently outperforms existing NLHF baselines, particularly excelling in scenarios involving diverse preferences and complex evaluation criteria.
Our analysis reveals that MNPO provides a unifying perspective on preference optimization, subsuming many existing methods as special cases while offering improved robustness in multi-agent alignment scenarios. By bridging recent advances in NLHF with the practical demands of aligning LLMs to diverse and potentially non-transitive human preferences, MNPO establishes a scalable foundation for next-generation alignment techniques.
2 RLHF Preliminaries
Notation.
$x\in\mathcal{X}$ denotes a prompt and $\mathcal{X}$ is the prompt space. $x$ is assumed to be sampled from a fixed but unknown distribution $d_{0}$. An LLM is characterized by a policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{Y})$ that takes a prompt $x$ as input and outputs a distribution over the response space $\mathcal{Y}$. The response $y\in\mathcal{Y}$ is then sampled from the distribution $\pi(\cdot\mid x)$.
Bradley-Terry Model Assumption.
The prevalent RLHF framework (Christiano et al., 2017; Ouyang et al., 2022) assumes the Bradley-Terry model. It assumes that there exists a reward function $r^{*}$ such that for any $x\in\mathcal{X}$ and $y^{1},y^{2}\in\mathcal{Y}$, we have $\mathbb{P}\left(y^{1}\succ y^{2}\mid x\right)=\frac{\exp\left(r^{*}\left(x,y^{1}\right)\right)}{\exp\left(r^{*}\left(x,y^{1}\right)\right)+\exp\left(r^{*}\left(x,y^{2}\right)\right)}=\sigma\left(r^{*}\left(x,y^{1}\right)-r^{*}\left(x,y^{2}\right)\right).$
After learning a reward function $R(\cdot,\cdot)$, RLHF algorithms aim to maximize the following KL-regularized objective with preference optimization RL algorithms such as PPO (Schulman et al., 2017):
$J(\pi)=\mathbb{E}_{x\sim d_{0}}\left[\mathbb{E}_{y\sim\pi(\cdot\mid x)}[R(x,y)]-\tau\operatorname{KL}\left(\pi(\cdot\mid x)\|\pi_{\text{ref}}(\cdot\mid x)\right)\right].$ (1)
Here $\pi_{\text{ref}}$ is the reference policy, which is usually a supervised fine-tuned LLM, and $\tau>0$ is the regularization parameter. By maximizing the objective, the obtained policy simultaneously achieves a high reward and stays close to $\pi_{\text{ref}}$, which can mitigate reward hacking (Tien et al., 2022; Skalse et al., 2022) to some extent.
General Preference Oracle.
The aforementioned algorithms rely on the Bradley-Terry model assumption, which may not always hold in practice. Recent studies (Munos et al., 2023; Calandriello et al., 2024; Ye et al., 2024; Zhang et al., 2025b; a) have instead directly considered the general preference distribution $\mathbb{P}$ without imposing additional assumptions, framing the preference optimization problem as a two-player game. These works assume the existence of a preference oracle $\mathbb{P}:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1]$. It can be queried to obtain binary preference signals as $z\sim\operatorname{Ber}\left(\mathbb{P}\left(y^{1}\succ y^{2}\mid x\right)\right)$, where $z=1$ indicates that $y^{1}$ is preferred to $y^{2}$, and $z=0$ indicates the opposite. The preference distribution is introduced as $\lambda_{\mathbb{P}}\left(x,y,y^{\prime}\right)=\left(y,y^{\prime}\right)\cdot\mathbb{I}\left[U<\mathbb{P}\left(y\succ y^{\prime}\mid x\right)\right]+\left(y^{\prime},y\right)\cdot\mathbb{I}\left[U\geq\mathbb{P}\left(y\succ y^{\prime}\mid x\right)\right]$, where $U\sim\operatorname{Uniform}(0,1)$ is a random variable.
Given two policies $\pi_{1}$ and $\pi_{2}$, LLMs are aligned using this general preference oracle. Consequently, the game objective is written as:
$J\left(\pi_{1},\pi_{2}\right)=\mathbb{E}_{x\sim D}[\mathbb{E}_{y_{1}\sim\pi_{1},y_{2}\sim\pi_{2}}\left[\mathbb{P}\left(y_{1}\succ y_{2}\mid x\right)\right]-\tau\operatorname{KL}\left(\pi_{1}\|\pi_{\text{ref}}\right)+\tau\operatorname{KL}\left(\pi_{2}\|\pi_{\mathrm{ref}}\right)]$ (2)
where the max-player $\pi_{1}$ aims to maximize the objective, and the min-player $\pi_{2}$ aims to minimize the objective. The goal of both players is to maximize their winning rates against the opponent while not deviating too far from $\pi_{\text{ref}}$, which shares a similar spirit with the objective of Eq. 1.
Nash Policy and Duality Gap.
Without loss of generality, we restrict our attention to the policy class $\Pi$ containing the policies with the same support set as $\pi_{\text{ref}}$. The Nash equilibrium of the game in Eq. 2 is then defined as:
$\pi_{1}^{*},\pi_{2}^{*}:=\underset{\pi_{1}\in\Pi}{\operatorname{argmax}}\,\underset{\pi_{2}\in\Pi}{\operatorname{argmin}}J\left(\pi_{1},\pi_{2}\right).$ (3)
Due to the symmetry of the two players, their Nash policies are unique and coincide, meaning $\pi_{1}^{*}=\pi_{2}^{*}=\pi^{*}$ (Ye et al., 2024). Interestingly, $J\left(\pi^{*},\pi\right)\geq 0.5$ always holds for $\forall\pi\in\Pi$, as $J\left(\pi^{*},\pi^{*}\right)=0.5$ indicates that $\pi^{*}$ is the best response against itself. To quantify how well a policy $\pi$ approximates the Nash policy $\pi^{*}$, we define the duality gap as:
$\operatorname{DualGap}(\pi):=\max_{\pi_{1}}J\left(\pi_{1},\pi\right)-\min_{\pi_{2}}J\left(\pi,\pi_{2}\right).$ (4)
The duality gap is nonnegative and $\operatorname{DualGap}(\pi)=0$ if and only if $\pi=\pi^{*}$. Hence, our goal is to find a policy that minimizes the duality gap. Once we achieve $\operatorname{DualGap}(\pi)\leq\epsilon$, we say that $\pi$ is an $\epsilon$-approximate Nash policy.
3 Method
3.1 RLHF as Multiplayer Games
To extend the two-player preference optimization objective to a multiplayer setting $\{\pi_{1},\pi_{2},...,\pi_{n}\}$, we consider a framework where each policy seeks to maximize its average preference probability against all other policies while regularizing toward a reference policy. We present the formulation of multiplayer games under two distinct assumptions: (i) the preference ranking assumption and (ii) the general preference oracle.
Plackett-Luce Reward Learning Assumption.
To generalize the maximum-likelihood reward learning objective for the Bradley-Terry model to one-vs-many comparisons, we adopt the Plackett-Luce framework. Specifically, we replace the pairwise logistic term $\log\sigma\left(R\left(x,y^{1}\right)-R\left(x,y^{2}\right)\right)$ with a softmax over multiple alternatives, extending the Bradley-Terry model to accommodate listwise comparisons. This Plackett-Luce model (Debreu, 1960; Plackett, 1975) maintains the interpretability of reward-based preferences while scaling to more complex decision-making scenarios. Formally, given a learned reward function $r(x,y)$ and a dataset $\mathcal{D}$ containing tuples $\left(x,\left\{y^{1},y^{2},\ldots,y^{k}\right\}\right)$, where $\left\{y^{1},y^{2},\ldots,y^{k}\right\}$ are a pool of $k$ to-be-ranked items, the probability that $y^{i}$ is preferred over the remaining pool of entities $\left\{y^{j}\right\}_{j\neq i}$ under the Plackett-Luce model is:
$\mathbb{P}\left(y^{i}\succ\left\{y^{j}\right\}_{j\neq i}\Bigm|x\right)=\frac{\exp({R\left(x,y^{i}\right)})}{\exp({R\left(x,y^{i}\right)})+\sum_{j\neq i}\exp({R\left(x,y^{j}\right)})}.$ (5)
The corresponding negative log-likelihood for a single comparison is:
$\log\mathbb{P}\left(y^{i}\succ\left\{y^{j}\right\}_{j\neq i}\Bigm|x\right)=\log\left(\exp\left(R\left(x,y^{i}\right)\right)+\sum_{j\neq i}\exp\left(R\left(x,y^{j}\right)\right)\right)-R\left(x,y^{i}\right).$ (6)
The generalized reward learning objective becomes:
[…]
Several key observations arise from this formulation. First, when $k=2$ for a one-vs-one comparison, Eq. 7 reduces to the vanilla Bradley-Terry objective, given by $\log\sigma\left(R\left(x,y\right)-R\left(x,y^{\prime}\right)\right)$, since $\sigma(a)=\frac{e^{a}}{1+e^{a}}$. Second, this objective maximizes the gap between the reward of $y$ and the log-sum-exp (LSE) of all alternatives, effectively applying a soft maximum over competitors. This penalizes cases where $y^{i}$ fails to dominate the collective "strength" of the dispreferred items $\left\{y^{j}\right\}_{j\neq i}$.
Multiplayer General Preference Optimization.
In addition to the Plackett-Luce assumption, we consider $n$-player preference optimization and leverage a universal preference oracle $\mathbb{P}:\mathcal{X}\times\mathcal{Y}\times\{\mathcal{Y}\}^{n-1}\rightarrow[0,1]$. It directly compares $y^{i}$ with a group of responses $\{y^{j}\}_{j\neq i}$, leading to the binary preference signals $z\sim\text{Ber}\left(\mathbb{P}\left(y^{i}\succ\left\{y^{j}\right\}_{j\neq i}\mid x\right)\right)$. Subsequently, each policy $\pi_{i}$ competes against other n1n-1 players, and the objective function becomes:
$J\left(\pi_{i},\left\{\pi_{j}\right\}_{j\neq i}\right)=\mathbb{E}_{x\sim D}\left[\mathbb{E}_{y^{i}\sim\pi_{i},\left\{y^{j}|y^{j}\sim\pi_{j}\right\}_{j\neq i}}\left[\mathbb{P}\left(y^{i}\succ\left\{y^{j}\right\}_{j\neq i}\Bigm|x\right)\right]-\tau\mathrm{KL}\left(\pi_{i}(\cdot\mid x)\|\pi_{\mathrm{ref}}(\cdot\mid x)\right)\right].$ (8)
Here, each policy $\pi_{i}$ seeks to maximize its expected preference over all other policies $\left\{\pi_{j}\right\}_{j\neq i}$. Meanwhile, the regularization term is kept to penalize $\pi_{i}$ with a KL divergence from the reference policy $\pi_{\text{ref}}$ weighted by $\tau$. This ensures that policies remain anchored to $\pi_{\text{ref}}$ to prevent overoptimization and maintain behavioral consistency. Regarding the dynamics of this multiplayer game, all policies $\left\{\pi_{i}\right\}_{i=1}^{n}$ are updated concurrently. Each policy $\pi_{i}$ optimizes its own $J$, leading to a competitive equilibrium where policies balance performance against the population and adherence to $\pi_{\mathrm{ref}}$.
Eq.8 has several key properties. (i) The symmetric treatment. All policies are treated equally, competing in a symmetric fashion. Therefore, $\pi_{1}^{*}=\pi_{2}^{*}=...=\pi_{n}^{*}$. (ii) Decentralized optimization. Each policy’s update depends only on its own actions and aggregate opponent behavior, avoiding complex interdependencies. (iii) Generalization of the two-player case. When $n=2$, each policy’s objective reduces to maximizing its pairwise preference probability minus its own KL penalty as in Eq. 2, aligning conceptually with competitive regularization.
Nash Equilibrium and Duality Gap.
In this $n$-player game, the Nash equilibrium is a policy $\pi^{*}$ where no player can improve their objective by unilaterally deviating. Formally, for all $\pi_{i}\in\Pi$ :
$J\left(\pi^{*}_{i},\{\pi^{*}_{j}\}_{j\neq i}\right)\geq J\left(\pi_{i},\{\pi^{*}_{j}\}_{j\neq i}\right),\quad\forall i\in\{1,\ldots,n\}.$ (9)
Due to symmetry, and assuming the KL terms vanish, the average win rate of the equilibrium policy $\pi^{*}$ against n1n-1 copies of itself is $\frac{1}{n}$. That is, $J\left(\pi^{*},\{\pi^{*}\}_{j\neq i}\right)=\frac{1}{n}$, analogous to the two-player case.
The duality gap quantifies how far a given policy $\pi$ is from the Nash policy $\pi^{*}$. Extending the two-player definition to an $n$-player setting, we define the set of opponent players as $O_{\pi}=\left\{\pi_{j}\right\}_{j=1}^{n-1}$. The duality gap in multiplayer games is then given by:
$\operatorname{DualGap}(\pi):=\mathbb{E}_{\pi_{j}\in O_{\pi}}\left[\max_{\pi_{j}}\min_{O_{\pi}\setminus\pi_{j}}J\left(\pi_{j},\pi\cup O_{\pi}\setminus\pi_{j}\right)\right]-\min_{O_{\pi}}J\left(\pi,O_{\pi}\right).$ (10)
This gap quantifies the maximum advantage a player could gain by unilaterally deviating from $\pi$ against the worst-case configuration of opponents, relative to the minimum payoff $\pi$ could incur if the opponents themselves deviate. Notably, the equilibrium condition is satisfied if and only if DualGap $\left(\pi^{*}\right)=0$, ensuring no player has an incentive to deviate unilaterally. Furthermore, if $\operatorname{DualGap}(\pi)\leq\epsilon$, then $\pi$ is defined as an $\epsilon$-approximate Nash policy in the multiplayer setting.
Multiplayer Nash Preference Optimization.
There are well-known algorithms that approximately solve the Nash equilibrium in a constant-sum multiplayer game. In this work, we follow (Freund & Schapire, 1999) to establish an iterative framework that can asymptotically converge to the optimal policy on average. Given a learning rate $\eta$ of online mirror descent update, we start with a theoretical analysis that conceptually solves the multiplayer game as follows  [^1] For notational conciseness and avoiding confusion, we use a superscript to denote time here, so that $\pi^{(t)}$ is equivalent to $\pi_{t}$. :
$\pi_{i}^{(t+1)}(y\mid x)\propto\left(\prod_{j\neq i}\pi_{j}^{(t)}(y\mid x)\right)^{\frac{1}{n-1}}\exp\left(\frac{\eta}{n-1}\sum_{j\neq i}\mathbb{P}\left(y\succ\pi_{j}^{(t)}\mid x\right)\right)$ (11)
Eq. 11 is an iterative framework that relies on the multiplicative weight update and enjoys a clear structure. Initially, we have a base policy $\pi^{(0)}$, usually from some supervised fine-tuned model $\pi_{\text{ref}}$. In each iteration $t$, the updated policy $\pi^{(t+1)}$ is obtained from the reference policy $\pi^{(t)}$ following the multiplicative weight update. Particularly, a response $y$ should have a higher probability weight if it has a higher average advantage over the current policy $\pi^{(t)}$. Equivalently, Eq. 11 can be written as $\pi^{(t+1)}_{i}(y\mid x)={\Pi_{j\neq i}{\pi^{(t)}_{j}(y\mid x)}^{\frac{1}{n-1}}\exp\left(\frac{\eta}{n-1}\mathbb{P}\left(y\succ\pi^{(t)}_{j}\mid x\right)\right)}/{Z_{\pi^{(t)}}({x})}$, where $Z_{\pi^{(t)}}({x})$ is the normalization factor (a.k.a, the partition function). Then for any fixed $x$ and $y$, each ideal update policy $\pi^{(t+1)}_{i}$ should satisfy:
$\frac{1}{n-1}\sum_{j\neq i}\log\frac{\pi^{(t+1)}_{i}(y\mid x)}{\pi^{(t)}_{j}(y\mid x)}=\frac{\eta}{n-1}\sum_{j\neq i}\mathbb{P}\left(y\succ\pi^{(t)}_{j}\mid x\right)-\log Z_{\pi^{(t)}}.$ (12)
Note that direct computation of $\pi^{(t+1)}$ involves a normalization factor, which is intractable for the exponentially large response space $\mathcal{Y}$. To avoid computing this normalization factor, we consider the logarithmic ratio between response pair $y$ and $y^{\prime}$, and define the function $h_{t}\left(\pi,y,y^{\prime}\right)$ as:
$h_{t}\left(\pi,y,y^{\prime}\right)=\log\frac{\pi(y\mid x)}{\pi(y^{\prime}\mid x)}-\frac{1}{n-1}\sum_{j\neq i}\left(\log\frac{\pi^{(t)}_{j}(y\mid x)}{\pi^{(t)}_{j}(y^{\prime}\mid x)}\right).$ (13)
From Eq.12, we know that the following equality holds for any response pair $y,y^{\prime}\in\operatorname{Supp}\left(\pi_{\text{ref}}\right)$:
$h_{t}\left(\pi^{(t+1)},y,y^{\prime}\right)=\frac{\eta}{n-1}\sum_{j\neq i}{\mathbb{P}\left(y\succ\pi^{(t)}_{j}\mid x\right)-\mathbb{P}\left(y^{\prime}\succ\pi^{(t)}_{j}\mid x\right)}$ (14)
Based on this observation, we define the loss function $L_{t}(\pi)$ and update the policy $\pi^{(t+1)}$ as:
$\pi^{(t+1)}\leftarrow\underset{\pi}{\operatorname{argmin}}\,\underbrace{\mathbb{E}_{y_{w},y_{l}\sim D_{t}}\left[\left(h_{t}\left(\pi,y_{w},y_{l}\right)-\frac{\eta}{n-1}\sum_{j\neq i}{\mathbb{P}\left(y\succ\pi^{(t)}_{j}\right)-\mathbb{P}\left(y^{\prime}\succ\pi^{(t)}_{j}\right)}\right)^{2}\right]}_{L_{t}(\pi)}$ (15)
It is clear to see that $\pi^{(t+1)}$ is the minimizer of $L_{t}(\pi)$ since $L_{t}\left(\pi^{(t+1)}\right)=0$. Furthermore, in the following lemma, we show that $\pi^{(t+1)}$ is the unique minimizer of $L_{t}$ within the policy class $\Pi$.
Lemma 1.
For each $t\in[T],\pi_{t+1}$ in Eq. 15 is the unique minimizer of $L_{t}(\pi)$ within $\Pi$.
The proof is deferred to Appendix E.1. Moreover, we replace the tricky term $\mathbb{P}\left(y\succ\pi^{(t)}_{j}\right)$ with a hyperparameter $\eta$ and propose the following loss to bypass it:
$L_{t}^{\prime}(\pi)=\mathbb{E}_{y,y^{\prime}\sim\pi_{t},\,y_{w},y_{l}\sim\lambda_{\mathbb{P}}\left(y,y^{\prime}\right)}\left[\left(h_{t}\left(\pi,y_{w},y_{l}\right)-\frac{1}{2\eta}\right)^{2}\right].$ (16)
Proposition 1.
For any policy $\pi\in\Pi$ and any iteration $t\in[T]$, $L^{\prime}_{t}(\pi)$ is equivalent to $L_{t}(\pi)$, differing only by an additive constant that is independent of $\pi$.
See the proof in Appendix E.2. Here, the response pair $(y,y^{\prime})$ is directly sampled from the current policy $\pi^{(t)}$, which is crucial for the equivalence between $L_{t}^{\prime}(\pi)$ and $L_{t}(\pi)$.
3.2 Multiplayer Nash Preference Optimization
Reward-Enhanced MNPO.
While our framework is motivated by moving beyond the limitations of purely scalar reward-based approaches, this does not preclude the beneficial incorporation of reward information when available. The key distinction lies in how rewards are utilized: rather than relying solely on reward-based rankings with implicit transitivity assumptions (as in classical RLHF), MNPO can leverage reward information as auxiliary guidance while maintaining the flexibility to handle non-transitive preferences through its game-theoretic structure.
Reward-aware preference optimization (RPO) (Sun et al., 2025) demonstrates how quantitative reward signals can complement qualitative preference comparisons. This approach aligns learned implicit preference models with explicit reward models that provide graded assessments of response quality. This shift allows MNPO to move beyond binary preference margins and instead minimize discrepancies between the learned reward function $r_{\pi_{\theta}}(x,y)$ and the target explicit reward model $r^{\star}(x,y)$. Concretely, instead of optimizing only the preference order between two responses, RPO defines the loss over preference pairs as:
$\mathcal{L}_{\text{RPO}}^{\mathbb{D}}\left(\pi_{\theta},\left(x,y^{1},y^{2}\right)\mid r^{\star},\pi_{\text{ref}},\beta,\eta\right)=:\mathbb{D}\left[r_{\pi_{\theta}}\left(x,y^{1}\right)-r_{\pi_{\theta}}\left(x,y^{2}\right)\|\eta r^{\star}\left(x,y^{1}\right)-\eta r^{\star}\left(x,y^{2}\right)\right],$ (17)
where $\left(x,y^{1},y^{2}\right)$ represents a preference pair and the distance metric $\mathbb{D}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}^{*}$ measures alignment between the model’s implicit reward differences and the scaled reference reward differences. The hyperparameters $\eta\in\mathbb{R}^{*}$ and $\beta\in\mathbb{R}^{*}$ control the reward scale and regularization, respectively. This formulation directly encourages the policy $\pi_{\theta}$ to internalize human-aligned reward values, bridging qualitative preference optimization with quantitative reward modeling.
[…]
Time-dependent Multiplayers
A central challenge in multiplayer optimization lies in defining and updating the set of opponent players $\left\{\pi_{j}^{(t)}\right\}_{j=1}^{n-1}$ in Eq.14. Inspired by recent iterative preference optimization methods such as DNO (Rosset et al., 2024), SPIN (Chen et al., 2024), and INPO (Zhang et al., 2025b), which typically rely on past policy iterations (e.g., $\pi_{\text{ref}}$ and $\pi^{(t-1)}$) to construct opponents, we adopt a time-dependent opponent selection mechanism.
Table 1: Time-dependent MNPO recovers many existing offline or online preference optimization algorithms. We denote the target reward gap as $\delta_{r^{\star}}:=\eta\left(r^{\star}\left(x,y_{1}\right)-r^{\star}\left(x,y_{2}\right)\right)$ . $\mathbb{D}^{\text{sq}}$ and $\mathbb{D}^{\text{bwd}}$ represent the squared distance and backward Bernoulli KL divergence, respectively.
Algorithm Num. Players Opponents Importance Weights Dist. Target Reward Gap
SimPO $n=1$ $\mathbb{D}^{\text{bwd}}$ $\infty$
DPO $n=2$ $\pi_{\text{ref}}$ $\lambda_{j}=1$ $\mathbb{D}^{\text{bwd}}$ $\infty$
Distill-DPO $n=2$ $\pi_{\text{ref}}$ $\lambda_{j}=1$ $\mathbb{D}^{\text{sq}}$ $\infty$
DNO $n=2$ $\pi_{t}$ $\lambda_{j}=1$ $\mathbb{D}^{\text{bwd}}$ $\infty$
SPIN $n=2$ $\pi_{t}$ $\lambda_{j}=\beta$ $\mathbb{D}^{\text{bwd}}$ $\infty$
SPPO $n=2$ $\pi_{t}$ $\lambda_{j}=1$ $\mathbb{D}^{\text{sq}}$ $\eta\left(\widehat{P}\left(y\succ\pi_{t}\mid x\right)-\frac{1}{2}\right)$
IPO $n=2$ $\pi_{\text{ref}}$ $\lambda_{j}=1$ $\mathbb{D}^{\text{sq}}$ $\frac{1}{2\tau}$
INPO $n=3$ $\pi_{t},\pi_{\text{ref}}$ $\lambda_{j}=\begin{cases}\frac{\tau}{\eta},&\text{if }j=t\\ \frac{\eta-\tau}{\eta},&\text{if }j=1\end{cases}$ $\mathbb{D}^{\text{sq}}$ $\frac{1}{2\tau}$
At any iteration $t$, we construct the opponent set from a mixture of recent historical policies $\{\pi_{t-j}\}_{j=0}^{t}$ ($n\leq t+1$), weighted by coefficients $\lambda_{j}$} with $\lambda_{j}\in[0,1]$ and optionally $\sum_{j}\lambda_{j}\leq 1$. The resulting time-dependent MNPO (TD-MNPO) loss is formulated as follows:
$\mathcal{L}_{\text{TD-MNPO}}^{t,\mathbb{D}}(\pi|\beta,\{\lambda_{j}\},\eta)=\mathbb{D}\left[\log\frac{\pi(y_{w}\mid x)}{\pi(y_{l}\mid x)}-\sum_{j=0}^{n-2}\lambda_{j}\log\frac{\pi_{t-j}(y_{w}\mid x)}{\pi_{t-j}(y_{l}\mid x)}\bigg\|\,\eta\delta^{\star}\right],$ (18)
where $\delta^{\star}$ encodes the target reward gap. By blending multiple past policies, this formulation stabilizes training, mitigates overfitting to transient fluctuations, and preserves temporal consistency.
Connections to Existing RLHF.
This unified MNPO formulation in Eq. 18 reveals that many preference optimization algorithms can be recovered as special cases by varying the number of players $n$, choice of opponents $O_{\pi}$, distance metric $\mathbb{D}$, and target reward gap $\delta^{\star}$.
For instance, DPO emerges by setting $n=2$, $O_{\pi}=\pi_{\text{ref}}$, and $\lambda_{j}=1$. Table 1 summarizes these reductions, showing how time-dependent MNPO unifies offline and online preference optimization under one principled framework. We provide a broader overview of RLHF objectives in Appendix D.
Compared to static reference-based approaches, the reward-aware multiplayer formulation offers: (i) Smoother policy evolution. Unlike methods that rely solely on the most recent past policy, MNPO gradually incorporates multiple past policies, preventing abrupt shifts and stabilizing policy updates. (ii) Greater robustness. By a weighted mixture of historical opponents, MNPO mitigates the risk of overfitting to transient fluctuations in recent iterations. (iii) Unified interpretation. TD-MNPO seamlessly extends existing approaches into a unified formulation, allowing for flexible adaptation to different training scenarios. (iv) Stable convergence.. The weighting scheme ensures that recent policies exert a stronger influence while preserving the broader learning trajectory, leading to more stable convergence. By dynamically leveraging historical policies as opponent players, TD-MNPO enhances preference optimization, making it more adaptable and robust in evolving learning environments. The pseudo-algorithm is described in Appendix 1.
4 Experimental Setup
Models and Training Settings.
We implement an online RLHF framework (Dong et al., 2024) with Gemma-2-9B-it (Team et al., 2024) as the base model. Our MNPO training consists of $T=3$ iterations, where each iteration generates responses from the current policy on a fresh prompt set and updates the policy using preference feedback. To eliminate the need for costly human annotations, we employ the reward model ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024a) to provide preference signals. Hyperparameter optimization proves critical, as optimal configurations vary both across different base models and between iterations of the same model. Through empirical analysis, we find that maintaining $\beta$ within the range $[0.01,10]$ consistently produces strong results. Furthermore, we observe that gradually increasing $\beta$ throughout training effectively mitigates training degradation while enabling continued model improvement. Complete implementation details and hyperparameter specifications are provided in Appendix C.
Evaluation Benchmarks.
We evaluate primarily on three widely used open-ended instruction-following benchmarks: MT-Bench (Zheng et al., 2023), AlpacaEval 2 (Li et al., 2023), and Arena-Hard v0.1 (Li et al., 2024). As suggested by Dubois et al. (2024), we report the win rate (WR) for Arena-Hard and the length-controlled (LC) WR for AlpacaEval 2, as judged by GPT-5-mini rather than the outdated GPT-4 Turbo (Preview-1106).
Since RLHF alignment is known to sometimes degrade reasoning, calibration, and factual accuracy (Ouyang et al., 2022; Dong et al., 2024), we further assess performance on a broader set of eleven academic benchmarks. These benchmarks span multiple abilities, including explicit instruction following (Zhou et al., 2023), general knowledge (Clark et al., 2018; Rein et al., 2024; Hendrycks et al., 2020), commonsense reasoning (Sakaguchi et al., 2021; Lin et al., 2021; Zellers et al., 2019), and math/coding problem solving (Lewkowycz et al., 2022; Chen et al., 2021).
In addition, we compare MNPO with a group of open-source LLMs, including LLaMA-3.1-8B-it (Dubey et al., 2024), Tulu-2-DPO-70B, LLaMA-3.3-70B-it, Mixtral-8x22B-it, and Qwen3-235B-it (Yang et al., 2025), and closed-source LLMs such as Gemini-2.5-Pro, GPT-5, and Claude-Sonnet-4.
Table 2: Performance of various models on instruction-following and preference-alignment benchmarks (AlpacaEval 2.0, Arena-Hard, and MT-Bench), evaluated using GPT-5-mini as the judge.
Model Size AlpacaEval 2.0 Arena-Hard MT-Bench
SFT Model 9B 50.15 44.97 6.49
DPO 9B 54.35 45.63 6.87
SimPO 9B 55.16 45.04 6.87
SPPO 9B 55.97 43.89 6.86
INPO 9B 56.09 48.03 6.95
MNPO 9B 57.27 52.26 7.03
LLaMA-3.1-8B-it 8B 5.24 39.16 5.37
Tulu-2-DPO-70B 70B 8.82 27.88 5.91
LLaMA-3.3-70B-it 70B 29.44 59.21 7.73
Mixtral-8x22B-it 141B 9.57 40.98 6.90
Qwen3-235B-it 235B 84.97 88.71 8.27
OpenAI/GPT-5 - 72.80 41.42 8.46
Anthropic/Claude-Sonnet-4 - 62.24 77.58 8.26
Google/Gemini-2.5-Pro - 90.93 86.98 7.78
Table 3: Model performance on instruction, knowledge, and commonsense benchmarks.
Model Instruction Knowledge Commonsense AVG
IFEval GPQA MMLU ARC HellaSwag TruthfulQA Winogrande
SFT Model 72.27 28.28 75.35 91.29 80.30 70.75 73.72 70.28
DPO 72.96 29.29 75.77 91.26 80.37 71.24 73.88 70.68
SimPO 73.79 32.32 76.79 91.09 78.90 63.40 72.93 69.60
SPPO 75.47 26.26 75.37 91.17 80.10 71.48 73.48 70.19
INPO 73.20 27.78 74.79 91.07 80.22 71.24 73.48 70.25
MNPO 73.94 33.33 75.63 91.15 80.18 70.26 73.09 71.08
Table 4: Model performance on math and coding benchmarks.
Model Math Code AVG
GSM8K Minerva-Math AIME-24 HumanEval
SFT Model 81.96 44.12 0 60.37 46.61
DPO 82.03 45.96 0 59.76 46.94
SimPO 82.56 43.38 0 57.32 45.82
SPPO 82.11 47.43 0 59.76 47.33
INPO 82.94 46.32 0 59.15 47.10
MNPO 82.64 44.85 3.33 61.59 48.10
5 Empirical Results
Instruction-Following and Preference Alignment.
Table 2 presents the performance of MNPO compared to existing preference optimization methods on three widely-used instruction-following benchmarks: AlpacaEval 2.0 (Length-Controlled Win Rate, %), Arena-Hard (Win Rate, %), and MT-Bench (Score/10). All models were evaluated using GPT-5-mini as the judge. MNPO consistently outperforms all baseline methods across all three benchmarks. On AlpacaEval 2.0, MNPO achieves a score of 57.27, representing improvements of 2.92 points over DPO (54.35), 2.11 points over SimPO (55.16), 1.30 points over SPPO (55.97), and 1.18 points over INPO (56.09). The improvements are even more pronounced on Arena-Hard, where MNPO scores 52.26 compared to the next best method INPO at 48.03, representing a 4.23-point improvement. Notably, on this challenging benchmark, MNPO not only excels over other preference optimization algorithms but also demonstrates strong competitiveness against much larger open-source fine-tuned models and even lastest closed-source models. It surpasses prominent models like Tulu2-DPO (70B) and Mixtral-it (141B), and, surprisingly, even outperforms GPT-5. On MT-Bench, MNPO achieves 7.03, outperforming all baselines, with the closest competitor being INPO at 6.95. These results demonstrate that the multiplayer formulation in MNPO provides significant advantages for instruction-following tasks. The consistent improvements across all three benchmarks suggest that the framework’s ability to handle diverse preferences and non-transitive relationships leads to better alignment with human expectations in open-ended generation tasks.
Knowledge and Reasoning Capabilities.
Table 3 evaluates model performance on academic benchmarks covering instruction following, knowledge, and commonsense reasoning. The results show that MNPO maintains strong performance across diverse cognitive tasks while achieving preference alignment. MNPO achieves the highest average score of 71.08 across all benchmarks, outperforming the SFT baseline (70.28) and all other preference optimization methods. Notably, MNPO achieves the best performance on GPQA (33.33), demonstrating strong graduate-level reasoning capabilities. The method also performs competitively on instruction-following (IFEval: 73.94) and maintains solid performance on knowledge benchmarks like MMLU (75.63) and commonsense reasoning tasks. Importantly, unlike some preference optimization methods that show degradation on certain academic benchmarks (e.g., SimPO’s drop to 63.40 on TruthfulQA), MNPO maintains relatively stable performance across all domains. This suggests that the multiplayer framework helps preserve the model’s foundational capabilities while improving preference alignment.
Mathematical and Coding Performance.
Table 4 presents results on mathematical reasoning and coding benchmarks. MNPO achieves the highest average score of 48.10 across math and coding tasks, outperforming all baseline methods, including SPPO (47.33) and INPO (47.10). On the challenging AIME-24 benchmark, MNPO is the only method to achieve non-zero performance (3.33), while all other methods including the SFT baseline score 0. This demonstrates MNPO’s superior capability in handling complex mathematical reasoning tasks. On HumanEval, MNPO achieves 61.59, representing the best coding performance among all methods. The results on GSM8K and Minerva-Math show that MNPO maintains competitive performance with existing methods while achieving superior results on the most challenging tasks. This pattern suggests that the multiplayer optimization framework is particularly beneficial for complex reasoning tasks that require handling multiple solution strategies or approaches.
6 Related Work
Reward-model–based RLHF.
Classical RLHF trains a reward model on human preference data and optimizes a policy with KL-regularized policy gradients (e.g., PPO) (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022; Schulman et al., 2017). While effective, PPO-style updates suffer from instability and high memory cost. To address this, GRPO removes the critic for more stable and memory-efficient training at scale (e.g., DeepSeek-R1) (Shao et al., 2024; Guo et al., 2025), and DAPO further improves sample efficiency through dynamic sampling and refined loss objectives (Yu et al., 2025). VAPO shows that value learning can strengthen RLHF when done properly (Yuan et al., 2025). Nonetheless, optimizing against imperfect reward proxies remains vulnerable to reward hacking (Weng, 2024; Wen et al., 2024).
RLHF with General Preference.
DPO bypasses reward modeling by directly optimizing a log-odds margin between preferred and dispreferred responses (Rafailov et al., 2023). This idea has inspired a family of extensions, including IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), SimPO (Meng et al., 2024), and WPO (Zhou et al., 2024), which refine the constraint, scaling, or sampling strategy. Iterative and online forms introduce exploration and continual updates (Xiong et al., 2023; Dong et al., 2024; Xie et al., 2024), but most remain limited to static pairwise supervision.
7 Conclusion
We introduce Multiplayer Nash Preference Optimization, extending Nash learning from human feedback to multiplayer settings. Our framework admits well-defined Nash equilibria and unifies existing preference optimization methods as special cases. Empirically, MNPO outperforms baselines across instruction-following and reasoning benchmarks. These results validate that multiplayer formulations better capture heterogeneous human preferences and provide more robust alignment for LLMs.
Appendix A Additional Related Work on Game-theoretic RLHF
Another perspective casts preference optimization as a game between the model and its opponents, where equilibrium-seeking methods provide stronger last-iterate guarantees. Self-play methods like SPIN (Chen et al., 2024), SPPO (Wu et al., 2024), and INPO (Zhang et al., 2025b) use no-regret dynamics, while Pre-DPO (Pan et al., 2025) and MPO (Wang et al., 2024b) adapt mirror descent. More recent methods, such as ONPO (Zhang et al., 2025a) and EGPO (Zhou et al., 2025), introduce optimism and extragradient techniques for stable convergence under noisy preferences. These advances primarily focus on two-player games but highlight the importance of equilibrium views in preference optimization. Extending RLHF to multiplayer interactions, as pursued by MNPO, generalizes beyond pairwise dynamics and opens the door to richer equilibrium structures for alignment.
Appendix B Pseudo-Algorithm of MNPO
Algorithm 1 Multiplayer Nash Preference Optimization (MNPO)
1:Require: Number of iterations $T$, distance metric $\mathbb{D}$, weight coefficients $\{\lambda_{j}\}$, reward scaling parameter $\eta$, regularization parameter $\beta$, external policy $\{\pi_{j}\}$, reference policy $\pi_{\text{ref}}$, policy class $\Pi$, preference oracle $\mathbb{P}$.
2:for iteration $t=1,2,\ldots,T$ do
3:  Use current policy $\pi_{t}$ to generate response pairs $\left\{y_{i}^{1},y_{i}^{2}\right\}_{i=1}^{m}$ where $y_{i}^{1},y_{i}^{2}\sim\pi_{t}$.
4:  Query the preference oracle $\mathbb{P}$ to get the preference dataset $D_{t}=\left\{y_{i}^{+},y_{i}^{-}\right\}_{i=1}^{m}$.
5:  Calculate $\pi_{t+1}$ as: $\pi_{t+1}\leftarrow\underset{\pi\in\Pi}{\operatorname{argmin}}\,\mathcal{L}_{\text{TD-MNPO}}^{t,\mathbb{D}}$
6:end for
7:Output $\pi_{T+1}$
Appendix C Experimental Details
Hardware and Implementation
All experiments are conducted on 8 NVIDIA H100 GPUs with 96GB memory. For the implemented algorithms, DPO (Rafailov et al., 2023) is trained using the official Hugging Face DPO Trainer, while SimPO (Meng et al., 2024) [^2] https://github.com/princeton-nlp/SimPO and SPPO (Wu et al., 2024) [^3] https://github.com/uclaml/SPPO follow their official GitHub implementations. INPO (Zhang et al., 2025b) is reproduced according to the settings described in the paper.
Hyperparameters
For MNPO, we adopt hyperparameters consistent with SimPO and INPO, using a cosine learning-rate scheduler with a peak learning rate of $5\times 10^{-7}$, a warmup ratio of 0.1, and a global batch size of 128. The optimizer is AdamW (Loshchilov & Hutter, 2017) without weight decay. We further perform a grid search for history weights at timesteps 1 and 2, selecting from $\{0,0.1,0.333,0.5,0.667,0.9\}$. In addition, we perform a grid search for $\eta$ over $\{0.1,0.01,0.0075,0.005,0.002\}$, and set $\eta=0.0075$.
Training Data
For training data, we use Gemma2-Ultrafeedback-Armorm (Cui et al., 2023) [^4] https://huggingface.co/datasets/princeton-nlp/gemma2-ultrafeedback-armorm, which contains approximately 60K training samples and 2K test samples. We retain both prompts and responses in iteration 1, while only prompts are used in iterations 2 and 3.
Evaluation Framework
For evaluation, we adopt the EvalScope framework (Team, 2024) [^5] https://github.com/modelscope/evalscope (version 1.0.2) across all datasets in Tables 3 and 4, and also applied it for AlpacaEval 2.0 and Arena-Hard in Table 2. MT-Bench [^6] https://github.com/lm-sys/FastChat/tree/main/fastchat/llm_judge is tested and implemented following its official GitHub repository. The LLM judge is configured with gpt5-mini-aug7-2025, where the reasoning effort is set to “minimal,” and all other parameters follow the default EvalScope settings.
Appendix D Formulations of Preference Optimization Objectives
Table 5 provides a consolidated overview of various preference optimization objectives that have been proposed in the literature. Each method is presented in terms of its optimization objective given preference data $\mathcal{D}=(x,y^{+},y^{-})$, where $x$ denotes the input prompt, and $y^{+}$ and $y^{-}$ denote the preferred (winning) and dispreferred (losing) responses, respectively. The table also specifies whether the method explicitly depends on a reference policy $\pi_{\text{ref}}$, whether it leverages the current policy $\pi_{t}$ during training, and whether the algorithm falls into the offline or online reinforcement learning regime.
Table 5: Various preference optimization objectives given preference data $\mathcal{D}=\left(x,y^{+},y^{-}\right)$, where $x$ is an input, and $y^{+}$ and $y^{-}$ are the winning and losing responses. $f$ is a class of divergence functions. $\Gamma(x,y)$ is the uncertainty estimator. $l$ is a convex decreasing loss function. $\widehat{P}\left(y\succ\pi_{t}\mid x\right)$ is the win rate over the distribution estimated by the average win rate over all the sampled responses $y_{1:K}\sim\pi_{t}(\cdot\mid x)$.
Method Objective $\pi_{\text{ref}}$ $\pi_{t}$ RL Type
DPO (Rafailov et al., 2023) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}-\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y^{+}\mid x\right)}{\pi_{\text{ref}}\left(y^{+}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y^{-}\mid x\right)}{\pi_{\text{ref}}\left(y^{-}\mid x\right)}\right)$ Offline
$f$-DPO (Wang et al., 2023) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}-\log\sigma\left(\beta f^{\prime}\left(\frac{\pi_{\theta}\left(y^{+}\mid x\right)}{\pi_{\mathrm{ref}}\left(y^{+}\mid x\right)}\right)-\beta f^{\prime}\left(\frac{\pi_{\theta}\left(y^{-}\mid x\right)}{\pi_{\mathrm{ref}}\left(y^{-}\mid x\right)}\right)\right)$ Offline
R-DPO (Park et al., 2024) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}-\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y^{+}\mid x\right)}{\pi_{\text{ref}}\left(y^{+}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y^{-}\mid x\right)}{\pi_{\text{ref}}\left(y^{-}\mid x\right)}+\left(\alpha\left|y^{+}\right|-\alpha\left|y^{-}\right|\right)\right)$ Offline
Distill-DPO (Fisch et al., 2024) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}\left[\log\frac{\pi_{\theta}\left(y^{+}\mid x\right)}{\pi_{\text{ref}}\left(y^{+}\mid x\right)}-\log\frac{\pi_{\theta}\left(y^{-}\mid x\right)}{\pi_{\text{ref}}\left(y^{-}\mid x\right)}-\left(r^{*}(x,y^{+})-r^{*}(x,y^{-})\right)\right]^{2}$ Offline
GSHF (Xiong et al., 2023) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}-\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y^{+}\mid x\right)}{\pi_{\text{ref}}\left(y^{+}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y^{-}\mid x\right)}{\pi_{\text{ref}}\left(y^{-}\mid x\right)}\right)+\left(\Gamma\left(x,y^{+}\right)-\Gamma\left(x,y^{-}\right)\right)$ Offline
KTO (Ethayarajh et al., 2024) $\begin{aligned} &\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}\left(y^{+}\mid x\right)}{\pi_{\text{ref}}\left(y^{+}\mid x\right)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}\left(y^{-}\mid x\right)}{\pi_{\text{ref}}\left(y^{-}\mid x\right)}\right),\\ &\text{ where }z_{\text{ref}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\beta\mathrm{KL}\left(\pi_{\theta}(y\mid x)\|\pi_{\text{ref}}(y\mid x)\right)\right]\end{aligned}$ Offline
IPO (Azar et al., 2024) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}\left[\log\frac{\pi_{\theta}\left(y^{+}\mid x\right)}{\pi_{\text{ref}}\left(y^{+}\mid x\right)}-\log\frac{\pi_{\theta}\left(y^{-}\mid x\right)}{\pi_{\text{ref}}\left(y^{-}\mid x\right)}-\frac{1}{2\tau}\right]^{2}$ Offline
SLiC-HF (Zhao et al., 2023) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}\max\left(0,\delta-\log\pi_{\theta}\left(y^{+}\mid x\right)+\log\pi_{\theta}\left(y^{-}\mid x\right)\right)-\lambda\log\pi_{\theta}\left(y^{+}\mid x\right)$ Offline
RRHF (Yuan et al., 2023) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}\max\left(0,-\frac{1}{\left|y^{+}\right|}\log\pi_{\theta}\left(y^{+}\mid x\right)+\frac{1}{\left|y^{-}\right|}\log\pi_{\theta}\left(y^{-}\mid x\right)\right)-\lambda\log\pi_{\theta}\left(y^{+}\mid x\right)$ Offline
SimPO (Meng et al., 2024) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}-\log\sigma\left(\frac{\beta}{\left|y^{+}\right|}\log\pi_{\theta}\left(y^{+}\mid x\right)-\frac{\beta}{\left|y^{-}\right|}\log\pi_{\theta}\left(y^{-}\mid x\right)-\gamma\right)$ Offline
CPO (Xu et al., 2024) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}-\log\sigma\left(\beta\log\pi_{\theta}\left(y^{+}\mid x\right)-\beta\log\pi_{\theta}\left(y^{-}\mid x\right)\right)-\lambda\log\pi_{\theta}\left(y^{+}\mid x\right)$ Offline
ORPO (Hong et al., 2024) $\begin{aligned} &\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}-\log p_{\theta}\left(y^{+}\mid x\right)-\lambda\log\sigma\left(\log\frac{p_{\theta}\left(y^{+}\mid x\right)}{1-p_{\theta}\left(y^{+}\mid x\right)}-\log\frac{p_{\theta}\left(y^{-}\mid x\right)}{1-p_{\theta}\left(y^{-}\mid x\right)}\right),\\ &\text{ where }p_{\theta}(y\mid x)=\exp\left(\frac{1}{|y|}\log\pi_{\theta}(y\mid x)\right)\end{aligned}$ Offline
DNO (Rosset et al., 2024) $\begin{aligned} \mathbb{E}_{\left(x,y^{\prime}_{t},y^{\prime\prime}_{t}\right)\sim\mathcal{D}_{t}}-\sigma\left(r_{t}\left(x,y^{\prime}_{t}\right)-r_{t}\left(x,y^{\prime\prime}_{t}\right)\right)\log\left[\sigma\left(\eta\log\frac{\pi\left(y^{\prime}_{t}\mid x\right)}{\pi_{t}\left(y^{\prime}_{t}\mid x\right)}-\eta\log\frac{\pi\left(y^{\prime\prime}_{t}\mid x\right)}{\pi_{t}\left(y^{\prime\prime}_{t}\mid x\right)}\right)\right]\\ -\sigma\left(r_{t}\left(x,y^{\prime\prime}_{t}\right)-r_{t}\left(x,y^{\prime}_{t}\right)\right)\log\left[\sigma\left(\eta\log\frac{\pi\left(y^{\prime\prime}_{t}\mid x\right)}{\pi_{t}\left(y^{\prime\prime}_{t}\mid x\right)}-\eta\log\frac{\pi\left(y^{\prime}_{t}\mid x\right)}{\pi_{t}\left(y^{\prime}_{t}\mid x\right)}\right)\right]\end{aligned}$ Online
DNO-Prct (Rosset et al., 2024) $\mathbb{E}_{\left(x,y_{t}^{+},y_{t}^{-}\right)\sim\mathcal{D}_{t}}-\log\left[\sigma\left(\widetilde{\eta}\log\frac{\pi\left(y_{t}^{+}\mid x\right)}{\pi_{t}\left(y_{t}^{+}\mid x\right)}-\widetilde{\eta}\log\frac{\pi\left(y_{t}^{-}\mid x\right)}{\pi_{t}\left(y_{t}^{-}\mid x\right)}\right)\right]$ Online
SPIN (Chen et al., 2024) $\mathbb{E}_{\left(x,y,y^{-}_{t}\right)\sim\mathcal{D}_{t}}-\ell\left(\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{t}(y\mid x)}-\beta\log\frac{\pi_{\theta}(y^{\prime}\mid x)}{\pi_{t}(y^{\prime}\mid x)}\right)$ Online
SPPO (Wu et al., 2024) $\mathbb{E}_{\left(x,y,\widehat{P}\left(y\succ\pi_{t}\mid x\right)\right)\sim\mathcal{D}_{t}}-\left[\log\frac{\pi_{\theta}(y\mid x)}{\pi_{t}(y\mid x)}-\eta\left(\widehat{P}\left(y\succ\pi_{t}\mid x\right)-\frac{1}{2}\right)\right]^{2}$ Online
INPO (Zhang et al., 2025b) $\mathbb{E}_{\left(x,y^{+}_{t},y^{-}_{t}\right)\sim D_{t}}-\left[\frac{\tau}{\eta}\left(\log\frac{\pi_{\theta}\left(y^{+}_{t}\right)}{\pi_{\mathrm{ref}}\left(y^{+}_{t}\right)}-\log\frac{\pi_{\theta}\left(y^{-}_{t}\right)}{\pi_{\mathrm{ref}}\left(y^{-}_{t}\right)}\right)+\frac{\eta-\tau}{\eta}\left(\log\frac{\pi_{\theta}\left(y^{+}_{t}\right)}{\pi_{t}\left(y^{+}_{t}\right)}-\log\frac{\pi_{\theta}\left(y^{-}_{t}\right)}{\pi_{t}\left(y^{-}_{t}\right)}\right)-\frac{1}{2\tau}\right]^{2}$ Online
ONPO (Zhang et al., 2025a) $\begin{aligned} &\mathbb{E}_{\left(x,y_{t}^{+},y_{t}^{-}\right)\sim D_{t}}\left[\log\frac{\pi_{\theta}\left(y^{+}_{t}\right)}{\pi_{t}^{\prime}\left(y^{+}_{t}\right)}-\log\frac{\pi_{\theta}\left(y^{-}_{t}\right)}{\pi_{t}^{\prime}\left(y^{-}_{t}\right)}-\frac{\eta}{2}\right]^{2},\\ &\text{ where }\pi_{t}^{\prime}=\underset{\pi}{\operatorname{argmin}}\mathbb{E}_{\left(x,y_{t}^{+},y_{t}^{-}\right)\sim D_{t}}\left[\log\frac{\pi\left(y^{+}_{t}\right)}{\pi_{t}\left(y^{+}_{t}\right)}-\log\frac{\pi\left(y^{-}_{t}\right)}{\pi_{t}\left(y^{-}_{t}\right)}-\frac{\eta}{2}\right]^{2}\end{aligned}$ Online
MIO (Lv et al., 2025) $\mathbb{E}_{\left(x,y^{+},y^{-}\right)\sim D}\left[\log\left(1+\frac{\pi_{\text{ref}}\left(y^{+}\mid x\right)}{\pi_{\theta}\left(y^{+}\mid x\right)}\right)+\frac{1}{2}\log\left(1+\frac{\pi_{\theta}\left(y^{+}\mid x\right)}{\pi_{\text{ref}}\left(y^{+}\mid x\right)}\right)+\frac{1}{2}\log\left(1+\frac{\pi_{\theta}\left(y^{-}\mid x\right)}{\pi_{\text{ref}}\left(y^{-}\mid x\right)}\right)\right]$ Offline
Appendix E Mathematical Analysis
E.1 Proof of Lemma 1
Proof.
First, by construction $L_{t}(\pi^{(t+1)})=0$, hence $\pi^{(t+1)}$ is a minimizer. Suppose, for contradiction, there exists $\tilde{\pi}\in\Pi$ with $\tilde{\pi}\neq\pi^{(t+1)}$ and $L_{t}(\tilde{\pi})=0$. Then for every pair $y,y^{\prime}\in\mathrm{Supp}(\pi_{\mathrm{ref}})$, we must have
$h_{t}(\tilde{\pi},y,y^{\prime})\;=\;\frac{\eta}{n-1}\sum_{j\neq i}\Bigl(\mathbb{P}(y\succ\pi^{(t)}_{j})-\mathbb{P}(y^{\prime}\succ\pi^{(t)}_{j})\Bigr)$.
Unrolling $h_{t}$ and rearranging yields the pairwise ratio identity
$\frac{\tilde{\pi}(y\,|\,x)}{\tilde{\pi}(y^{\prime}\,|\,x)}=\frac{\exp\!\Bigl(\tfrac{\eta}{n-1}\sum_{j\neq i}\mathbb{P}(y\succ\pi^{(t)}_{j}\,|\,x)\Bigr)\prod_{j\neq i}\pi^{(t)}_{j}(y\,|\,x)^{\frac{1}{n-1}}}{\exp\!\Bigl(\tfrac{\eta}{n-1}\sum_{j\neq i}\mathbb{P}(y^{\prime}\succ\pi^{(t)}_{j}\,|\,x)\Bigr)\prod_{j\neq i}\pi^{(t)}_{j}(y^{\prime}\,|\,x)^{\frac{1}{n-1}}}.$
Because $\mathrm{Supp}(\tilde{\pi})=\mathrm{Supp}(\pi_{\mathrm{ref}})$ and $\sum_{y}\tilde{\pi}(y\,|\,x)=1$, these pairwise ratios determine $\tilde{\pi}(\cdot\,|\,x)$ uniquely—and that unique solution is exactly the normalized distribution implied by Eq. 11 and 12, i.e., $\pi^{(t+1)}$. Hence $\tilde{\pi}=\pi^{(t+1)}$, a contradiction. Therefore the minimizer is unique. ∎
E.2 Proof of Proposition 1
Proof.
Introduce an indicator $I\sim\mathrm{Ber}(\mathbb{P}(y\succ y^{\prime}\,|\,x))$ for a pair $(y,y^{\prime})\sim\pi^{(t)}\times\pi^{(t)}$. Consider
$\widetilde{L}_{t}(\pi)\;=\;\mathbb{E}_{y,y^{\prime}\sim\pi^{(t)},\;I}\Bigl(h_{t}(\pi,y,y^{\prime})-\tfrac{I}{\eta}\Bigr)^{2}.$
Expanding and comparing $\widetilde{L}_{t}(\pi)$ with $L_{t}(\pi)$ shows the only difference lies in the cross-term $\mathbb{E}[\,h_{t}(\pi,y,y^{\prime})(\mathbb{P}(y\succ\pi^{(t)})-\mathbb{P}(y^{\prime}\succ\pi^{(t)}))\,]$ versus $\mathbb{E}[\,h_{t}(\pi,y,y^{\prime})I\,]$. One verifies these are equal by writing $h_{t}$ as a linear form in $\log\pi$, $\log\pi^{(t)}_{j}$ and using that $y$ and $y^{\prime}$ are i.i.d. draws from $\pi^{(t)}$:
$\mathbb{E}_{y,y^{\prime}}\!\big[h_{t}(\pi,y,y^{\prime})\big(\mathbb{P}(y\succ\pi^{(t)})-\mathbb{P}(y^{\prime}\succ\pi^{(t)})\big)\big]=\mathbb{E}_{y,y^{\prime},I}\!\big[h_{t}(\pi,y,y^{\prime})\,I\big].$
(See INPO Appendix A.5 for the algebraic steps; the same symmetry argument applies verbatim.) Now expand $L^{\prime}_{t}(\pi)$ by conditioning on $(y,y^{\prime})$ and the preference sampler $\lambda_{\mathbb{P}}(y,y^{\prime})$:
$L^{\prime}_{t}(\pi)=\mathbb{E}_{y,y^{\prime}}\Bigl[\mathbb{P}(y\succ y^{\prime})\Bigl(h_{t}(\pi,y,y^{\prime})-\tfrac{1}{2\eta}\Bigr)^{2}+\bigl(1-\mathbb{P}(y\succ y^{\prime})\bigr)\Bigl(h_{t}(\pi,y^{\prime},y)-\tfrac{1}{2\eta}\Bigr)^{2}\Bigr].$
Using $h_{t}(\pi,y^{\prime},y)=-h_{t}(\pi,y,y^{\prime})$ and completing the square shows $L^{\prime}_{t}(\pi)=\widetilde{L}_{t}(\pi)+\mathrm{const}$, where the constant depends only on the distribution of $(y,y^{\prime})$ and $\mathbb{P}(\cdot)$, not on $\pi$. Combining with the equality of cross-terms above gives $L^{\prime}_{t}(\pi)=L_{t}(\pi)+\mathrm{const}$, as claimed. ∎
E.3 Theoretical Analysis of Eq. 11
Here, we provide a theoretical analysis of the iterative update in Eq.  11:
$\pi^{(t+1)}_{i}(y\mid x)\;\propto\;\Biggl(\prod_{j\neq i}\pi^{(t)}_{j}(y\mid x)\Biggr)^{\!\frac{1}{n-1}}\exp\!\Biggl(\frac{\eta}{n-1}\sum_{j\neq i}P\!\left(y\succ\pi^{(t)}_{j}\,\middle|\,x\right)\Biggr).$ (19)
Derivation from Mirror Descent.
Eq. 19 can be viewed as an instance of online mirror descent (OMD) with the KL divergence as the Bregman potential. At each step, player $i$ seeks to maximize the expected win probability against the population $\left\{\pi^{(t)}_{j}\right\}_{j\neq i}$ subject to a KL regularization toward the opponent mixture:
$\pi^{(t+1)}_{i}=\operatorname*{arg\,max}_{\pi\in\Pi}\;\frac{1}{n-1}\sum_{j\neq i}\Bigl\langle\pi,\,P(\cdot\succ\pi^{(t)}_{j})\Bigr\rangle-\frac{1}{\eta}\,D_{\mathrm{KL}}\!\left(\pi\,\middle\|\,\Bigl(\!\prod_{j\neq i}\pi^{(t)}_{j}\Bigr)^{\!\frac{1}{n-1}}\right).$
Taking first-order conditions yields exactly the multiplicative-weights update in Eq. 19. Thus, the MNPO update inherits the regret guarantees of OMD: the average regret after $T$ rounds scales as $\mathcal{O}(1/\sqrt{T})$, ensuring convergence to equilibrium in the no-regret learning sense.
Pairwise Ratio Dynamics.
To avoid computing the intractable partition function, we analyze the pairwise log-ratio
$h_{t}(\pi,y,y^{\prime})\;=\;\log\frac{\pi(y\mid x)}{\pi(y^{\prime}\mid x)}-\frac{1}{n-1}\sum_{j\neq i}\log\frac{\pi^{(t)}_{j}(y\mid x)}{\pi^{(t)}_{j}(y^{\prime}\mid x)}.$
Eq. 12 in the main text shows that at the fixed point $\pi^{(t+1)}$, these ratios satisfy
$h_{t}\!\left(\pi^{(t+1)},y,y^{\prime}\right)\;=\;\frac{\eta}{n-1}\sum_{j\neq i}\Bigl(P(y\succ\pi^{(t)}_{j}\mid x)-P(y^{\prime}\succ\pi^{(t)}_{j}\mid x)\Bigr).$
Hence the log-ratio dynamics correspond to a consistent linearization of the preference margins across all opponents, ensuring that the update increases probability mass on responses with strictly higher average advantage.
Equilibrium Properties.
By standard arguments in no-regret game dynamics (Freund & Schapire, 1999), if all players update according to Eq. 19, the joint empirical distribution converges to an $n$-player Nash equilibrium. Moreover, because the update is multiplicative in form, probabilities remain strictly positive on the support of $\pi_{\mathrm{ref}}$, preventing premature collapse.
Interpretation.
Eq. 19 admits two complementary interpretations:
  • Population averaging: the geometric mean term aggregates beliefs of all opponents, ensuring stability against heterogeneous policies.
  • Advantage weighting: the exponential term amplifies responses that consistently outperform others, with learning rate $\eta$ controlling the exploration–exploitation trade-off.
Together, these properties explain why MNPO achieves robust convergence in multiplayer preference optimization while generalizing the two-player INPO update.
Appendix F Limitations and Future Work
Like other algorithms in the RLHF paradigm, the performance of MNPO is fundamentally linked to the quality of its preference data. Two primary limitations warrant consideration for future work.
First, the fidelity of the preference oracle serves as a performance ceiling. As the policy model improves and its generations become consistently high-quality, the task of distinguishing between chosen and rejected responses becomes increasingly difficult for the preference oracle in practice. This diminishing discriminative capability can become a bottleneck for further improvement.
Second, the paradigm of simply increasing the probability of chosen responses and decreasing that of rejected ones may face diminishing returns as the preference gap narrows. When rejected responses are themselves of high quality, the binary preference signal becomes less informative, potentially slowing down or stalling the convergence of the policy model. Future research could explore more nuanced feedback mechanisms to address learning in this high-performance regime.
F.1 External Opponent Players
Formulation.
An alternative to using past-time policies as opponents is leveraging external LLMs to introduce a broader range of competitive dynamics. Given a set of n1n-1 LLM policies $\{\pi_{j}\}_{j=1}^{n-1}$, these opponent models can come from different sources: they may belong to distinct model families, have varying parameter scales within the same architecture, be trained on different data distributions, or specialize in different domains. The external opponent MNPO (EO-MNPO) loss in this case is defined as:
$\mathcal{L}_{\text{EO-MNPO}}^{t,\mathbb{D}}(\pi|\beta,\{\lambda_{j}\},\eta)=\mathbb{D}\left[\sum_{j}\lambda_{j}\left(\log\frac{\pi(y\mid x)}{\pi_{j}(y\mid x)}-\log\frac{\pi(y^{\prime}\mid x)}{\pi_{j}(y^{\prime}\mid x)}\right)\bigg\|\eta\delta^{\star}\right].$ (20)
Here, the constraint $\sum_{j}\lambda_{j}=1$ ensures that the contributions of different opponent policies are appropriately weighted. This formulation introduces a key advantage: it enables preference optimization across diverse knowledge sources rather than being restricted to a single training trajectory.
To better interpret this formulation, we can draw parallels to knowledge distillation. Specifically, consider a scenario where we have a collection of teacher models $\left\{\pi^{\text{teacher}}_{i}\right\}$ and seek to refine an updated policy $\pi$ that remains close to both these teacher models and the original reference model $\pi_{\text{ref}}$. The objective function can be expressed as:
$J(\pi)=\mathbb{E}_{x\sim d_{0}}\left[\mathbb{E}_{y\sim\pi}[R(x,y)]-\tau_{0}\operatorname{KL}\left(\pi\|\pi_{\text{ref}}\right)-\sum_{i}\tau_{i}\operatorname{KL}\left(\pi\|\pi^{\text{teacher}}_{i}\right)\right].$ (21)
This formulation illustrates that MNPO with external opponent players can be viewed as an extension of knowledge distillation, where the learned policy integrates information from multiple expert models while balancing divergence from the reference model. Using the same derivation technique as DPO, we can recover Eq. 20, linking preference optimization to a generalized knowledge alignment framework.
Proposition 2.
The optimal solution to the maximum reward objective in Eq. 21 is equivalent to the learned reward model in Eq. 20.
It is straightforward to show that the solution to the maximum reward objective in Eq. 21 takes the form: $\pi^{\star}(y\mid x)=\frac{1}{Z(x)}{\exp(R(x,y)/\tau)\pi_{\mathrm{ref}}(y\mid x)^{\tau_{0}/\tau}\prod_{i}\pi_{i}^{\text{teacher }}(y\mid x)^{\tau_{i}/\tau}}$, where $Z(x)$ is the partition function and $\tau=\tau_{0}+\sum_{i}\tau_{i}$.
Analysis.
EO-MNPO gains several benefits. Firstly, Unlike the self-comparison TD-MNPO, using a diverse pool of external LLMs exposes the policy to a broader range of feedback, leading to more robust optimization. Secondly, leveraging domain-specific expert models allows MNPO to fine-tune policies for specialized applications, improving generalization. Thirdly, this framework aligns with broader trends in multi-agent RL, where agents iteratively refine their strategies against multiple opponents, driving more sophisticated decision-making. By considering a dynamic set of external LLMs as opponent players, EO-MNPO extends beyond self-referential optimization, making it a more flexible and generalizable approach for preference optimization in LLMs.
Appendix G The Use of Large Language Models (LLMs)
We used large language models, including ChatGPT, Gemini, and Claude as writing assistants in the preparation of this manuscript. Their role was strictly confined to language enhancement tasks, such as polishing grammar, improving readability, and generating alternative phrasings. All intellectual contributions, such as research ideas, experimental designs, analyses, and conclusions, were developed solely by the authors, who take full responsibility for the content of this paper.